1. Which term of the AP: 121, 117, 113, . . ., is its first negative term? [Hint: Find n for an < 0]

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Given the AP series is 121, 117, 113, . . ., Thus, first term, a = 121 Common difference, d = 117-121= -4 By the nth term formula, an = a+(n −1)d Therefore, an = 121+(n−1)(-4) = 121-4n+4 =125-4n To find the first negative term of the series, an < 0 Therefore, 125-4n < 0 125 < 4n n>125/4 n>31.25 Therefore, the first negative term of the series is 32nd term.